Ultravision Inc. anticipates sales of $300,000 from January through April. Materials will represent 50 percent of sales, and because of level production, material purchases will be equal for each month during the four months of January, February, March, and April.

Materials are paid for one month after the month purchased. Materials purchased in December of last year were $26,000 (half of $52,000 in sales). Labor costs for each of the four months are slightly different due to a provision in the labor contract in which bonuses are paid in February and April. The labor figures are:

January $16,000
February 19,000
March 16,000
April 21,000

Fixed overhead is $12,000 per month.

Prepare a schedule of cash payments for January through April. (Assume the $300,000 of sales occur equally over the four months of January through April, i.e. Monthly sales = $300,000 / 4.)
Answer #1
Ultravision Inc
schedule of cash payments  
For the months of January to April
January February March   April
Cash paid for purchases $   26,000 $300,000/4*50% = $37,500 $300,000/4*50% = $37,500 $300,000/4*50% = $37,500
Cash for salaries $   16,000 $                                       19,000 $                                       16,000 $                                       21,000
Cash paid for fixed overhead $   12,000 $                                       12,000 $                                       12,000 $                                       12,000
Total cash payments $   54,000 $                                       68,500 $                                       65,500 $                                       70,500

You can reach me over comment box if you have any doubts. Please rate this answer